OpenStudy (anonymous):

Simplify: (5+2v7)^2 A. 25+4v7 B. 53+20v7 C. 53

OpenStudy (anonymous):

I don't think you have the problem well represented. Try again, and maybe we can figure it out.

OpenStudy (anonymous):

i just edited it

OpenStudy (anonymous):

2v7? Not sure what that means.

OpenStudy (anonymous):

Sqrt .

OpenStudy (anonymous):

I got A. for answer

OpenStudy (anonymous):

Correct ?

OpenStudy (anonymous):

I don't think so... give me a second.\[(5+2\sqrt{7})^2 =25+20\sqrt{7}+28=53+20\sqrt{7}\]

OpenStudy (anonymous):

You've got to be disciplined about using the FOIL method on this kind of problem, or you can use the pattern for sum of squares\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

Got it now?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Good.